# Ting En Lam1,2, Yuhan Chen1,3, Elston Tan1,4, Eric Peh1, Ruirui Chen1, Paritosh Parmar1, Basura Fernando1

### v1

@@ 4.2 Benchmarking on OEAG
In the next step, we integrated visual information into LLM. Taking inspiration from [31], we learn a projector network to align the visual features with GPT-2 representations (VisualGPT-2 in Table 3(a)). We observed an improvement in the performance. However, we believe that this might not be a very efficient way to integrate visual information with LLMs. We expect the performance to boost considerably through better and more sophisticated joint modeling vision and language. Overall, we believe shallower networks might not have the capacity to do inference over longer, complex causal chains, and as such it might not be the best option to invest future efforts into; LLMs have the potential to excel at causal reasoning; this view is also supported in a concurrent survey [25].
@@
Table 4: Our dataset improves performance on existing real-world dataset. For OEAG, CapsMIX score is reported.
@@ 4.3 Does our dataset help with real-world cases?
To observe our dataset’s direct and immediate impact, we combine our training set with NextQA’s training set in this experiment. Then, we note the performance of models on NextQA test set. For comparison, we also note the performance of models when our dataset is not leveraged/incorporated. Despite being a synthetic/cartoon dataset, our dataset boosted the performance on a real-world dataset on both MCQA and OEAG (Table 4). We observed improvement in identifying the correct causes by breaking the reliance on shortcuts and focus on causal effects; a more comprehensive analysis of the situation/interaction, rather than jumping to a conclusion. Qualitative results provided in Supplementary material. What is more, improvements were not limited to ‘causal-why’ questions but also extended across other question types. We hypothesize that this can attributed to the wider range of reasoning involved in our dataset. To the best of our knowledge, this is the first time a synthetic VideoQA dataset has shown immediate improvement on a real-world dataset. However, we noticed a slight drop in the performance on location-type questions. Thus, we believe that it might be better to: 1) transfer reasoning skills acquired from challenging synthetic datasets like ours; 2) leverage our synthetic dataset to inform the model designing process, as it better reflects the challenges VideoQA models may face in the real world, such as longer causal chains and frequent scene changes in visual streams. Nonetheless, we do not recommend naively transferring or deploying whole models/weights trained on our dataset to real-world scenarios or applications.
@@ 5 Conclusion
We introduced CausalChaos!, a challenging dataset for causal action question-answering tasks. Based on the timeless classic of the Tom and Jerry cartoon, it is richly annotated with critical thinking questions demanding a wide spectrum of reasoning from Video QA models. Questions are supported with multi-level answers and further explanations, all-encompassing video context. We further provide a novel CausalConfusion test set, which can challenge the causal relationship modeling capabilities of Video QA models. Through experimentation, we found that while existing models work well on causal action QA tasks, there is a lot of room to improve, especially, in terms of causal relationship modeling, and long and detailed open-ended answer generation. While LLMs offer a promising start, sophisticated integration of visual information with LLMs or joint modeling vision and language would be crucial. We hope that our challenging dataset will foster such developments. We will release our dataset and codes to help future efforts in this direction. Lastly, we showed that it brings improvement to real-world datasets. We believe our dataset, derived from cartoons, should inform the model designing process, as it better reflects the challenges Video-QA models may face in the real world, such as longer causal chains and frequent scene changes in visual streams. However, we do not suggest naively transferring or deploying whole models trained on our dataset to real-world scenarios.

### v2

@@ 4.2 Benchmarking on Open-ended Answer Generation (OEAG)
In the next step, we integrated visual information into LLM. Taking inspiration from [30], we learn a projector network to align the visual features with GPT-2 representations (VisualGPT-2 in Table 3. We observed an improvement in the performance. However, we believe that this might not be a very efficient way to integrate visual information with LLMs. We expect the performance to boost considerably through better and more sophisticated joint modeling vision and language. Overall, we believe shallower networks might not have the capacity to do inference over longer, complex causal chains, and as such it might not be the best option to invest future efforts into; LLMs have the potential to excel at causal reasoning; this view is also supported in a concurrent survey [25].
@@
Table 4: Our dataset improves performance on existing real-world dataset. For OEAG, WUPS score is reported.
@@ 4.3 Does our dataset help with real-world cases?
To evaluate the direct impact of our dataset, we combine our training set with NextQA’s training set and measure model performance on the NextQA test set. For comparison, we also measure model performance without incorporating our dataset. Despite being a synthetic/cartoon dataset, and much smaller in size, our dataset boosted the performance on a real-world dataset on both MCQA and OEAG (Table 4). We observed improvement in identifying the correct causes by breaking the reliance on shortcuts and focus on causal effects; a more comprehensive analysis of the situation/interaction, rather than jumping to a conclusion. Qualitative results provided in Appendix. What is more, improvements were not limited to ‘causal-why’ questions but also extended across other question types. We hypothesize that this can attributed to the wider range of reasoning involved in our dataset. To the best of our knowledge, this is the first time a synthetic VideoQA dataset has shown immediate improvement on a real-world dataset. However, we noticed a slight drop in the performance on location-type questions. Thus, we believe that it might be better to: 1) transfer reasoning skills acquired from challenging synthetic datasets like ours; 2) leverage our synthetic dataset to inform the model designing process, as it better reflects the challenges VideoQA models may face in the real world, such as longer causal chains and frequent scene changes in visual streams. Nonetheless, we do not suggest naively deploying whole models/weights trained on our dataset to real-world scenarios or applications.
@@ 5 Conclusion
We introduced CausalChaos!, a challenging dataset for causal action question-answering tasks based on the classic Tom and Jerry cartoon, richly annotated with critical thinking questions requiring extensive reasoning from Video QA models. Questions come with multi-level answers and explanations covering the entire video context. We also provide the novel CausalConfusion test set to challenge causal relationship modeling in Video QA models. Our experiments show that while existing models perform well on causal action QA tasks, there is significant room for improvement in causal relationship modeling and generating detailed open-ended answers. LLMs show promise, but integrating visual information with LLMs or joint modeling of vision and language is crucial. We hope our dataset fosters such developments and will release it and the codes to support future efforts. Lastly, we demonstrated improvements in real-world datasets. Our dataset, derived from cartoons, should inform model design, reflecting real-world challenges like longer causal chains and frequent scene changes. However, we do not suggest deploying models trained solely on our dataset in real-world scenarios.
Material Acknowledgement and Disclaimer.
Tom and Jerry is a material of Turner Entertainment Company (Warner Bros. Entertainment Inc.). All rights reserved. We do not claim any ownership of or rights to the Tom and Jerry material. All other trademarks, service marks, trade names and any other material referenced in this document are the property of their respective owners.